3^(x+3)=5^x
thank you

Answer by stanbon(75887)

Take the log of both sides to get:
(x+5)log3 = xlog5
(x+5)/x = log5/log3=1.465
1+(5/x)=1.465
5/x =0.465
x/5= 2151
x=10.753
Cheers,
Stan H.